# Sublinear Data Structures for Nearest Neighbor in Ultra High Dimensions

### v1

@@ 3 Preliminaries
$\underset{}{\mathbb{P}}\left[X\geq(1+\delta)\mu\right]\leq\exp(-\delta^{2}\mu/(3c)),\text{ for }0<\delta\leq 1$
$\underset{}{\mathbb{P}}\left[X\geq(1+\delta)\mu\right]\leq\exp(-\delta\mu/(3c)),\text{ for }\delta\geq 1$
$\underset{}{\mathbb{P}}\left[X\leq(1-\delta)\mu\right]\leq\exp(-\delta^{2}\mu/(2c)),\text{ for }\delta\geq 0.$
@@
4 Warm-Up: Data Structure for $\ell_{1}$ Metric with $\tilde{O}(n^{2})$ Space and $\tilde{O}(n)$ Query Complexity
@@
Definition 4.1 (Scaled Bounding Box).
For any two points $a,b\in\mathbb{R}^{d}$, the scaled bounding box $B_{a,b}$ specified by $(a,b)$ is $\left\{x\in\mathbb{R}^{d}\mid\forall z\in[d],\min\{a^{(z)},b^{(z)}\}-100\cdot|a^{(z)}-b^{(z)}|\leq x^{(z)}\leq\max\{a^{(z)},b^{(z)}\}+100\cdot|a^{(z)}-b^{(z)}|\right\}$.
@@
Theorem 4.2 is a special case of Theorem 6.1, where a rigorous proof is provided. In this section, we present a more concise proof to provide intuition. Our approach is the following. We first describe a data structure for the approximate nearest neighbor problem for two points (i.e. a 2-point comparator), and then employ a naive min-finding procedure with oracle access to the 2-point comparator, to extend to the approximate nearest neighbor with $n$ points.
4.1 A Strong Comparator for Two Points
@@
We analyze the correctness of Algorithm 1 in Lemma 4.3. In particular, we show the following strong guarantees: (i) if one point is significantly closer to the query than the other point, the comparator outputs the correct point with probability $1-\delta$, (ii) otherwise, the comparator may output a special symbol $\perp$, but never outputs a wrong point as the nearest neighbor.
@@ 10       Let $x,y\leftarrow 0$;
Lemma 4.3 (Strong Comparator Correctness).
@@
Since $\mathbb{E}[Y_{t}]=\|b-q\|_{1}\geq\frac{1}{2}\|b-a\|_{1}$, using Chernoff bound, we have $\Pr\left[|y-\mathbb{E}[y]|>\frac{\epsilon}{8}\mathbb{E}[y]\right]\leq\frac{\delta}{2}$ by choosing $T=\Theta(\log(1/\delta)/\epsilon^{2})$. To analyze $x$, we consider the following two cases:
Case 1. If $\mathbb{E}[X_{t}]=\|a-q\|_{1}\geq\frac{1}{8}\|b-a\|_{1}$, then using Chernoff bound, we have $\Pr\left[|x-\mathbb{E}[x]|>\frac{\epsilon}{8}\mathbb{E}[x]\right]\leq\frac{\delta}{2}$ by choosing $T=\Theta(\log(1/\delta)/\epsilon^{2})$.
Case 2. If $\mathbb{E}[X_{t}]=\|a-q\|_{1}<\frac{1}{8}\|b-a\|_{1}$, then using Chernoff bound, we have $\Pr\left[|x-\mathbb{E}[x]|>\frac{T}{8}\|b-a\|_{1}\right]\leq\frac{\delta}{2}$ by choosing $T=\Omega(\log(1/\delta))$.
To prove item 1 of Lemma 4.3, suppose that $\frac{\|b-q\|_{1}}{\|a-q\|_{1}}\geq 1+\epsilon$. If we are in Case 1, with probability $1-\delta$ we have $x\approx_{\frac{\epsilon}{8}}T\|a-q\|_{1}$ and $y\approx_{\frac{\epsilon}{8}}T\|b-q\|_{1}$. This implies that, with probability $1-\delta$ we get $\frac{y}{x}\geq\frac{1-\frac{\epsilon}{8}}{1+\frac{\epsilon}{8}}\cdot(1+\epsilon)\geq 1+\frac{\epsilon}{2}$. Thus the querying phase correctly outputs $a$ as the nearest neighbor. If we are in Case 2, with probability $1-\delta$ we have $x\leq\frac{1}{4}T\|b-a\|_{1}$, and $y\approx_{\frac{\epsilon}{8}}T\|b-q\|_{1}$, which implies that $\frac{y}{x}\geq\frac{(1-\frac{\epsilon}{8})(1-\frac{1}{8})}{\frac{1}{4}}\geq 3$. Thus the querying phase correctly outputs $a$.
The second item of Lemma 4.3 can be proved similarly. ∎
4.2 Data Structure for $\ell_{1}$ Metric with $\tilde{O}(n^{2})$ Space
@@ 10                   Let $j\leftarrow i$;
Proof.
(for Theorem 4.2) By a union bound, we get that with probability $1-\delta$, for all $1\leq i<j\leq n$, the data structure $A_{(c_{i},c_{j})}$ does not fail. We condition on this event in the rest of the proof. Notice that $j$ is a variable under updating.
Algorithm 2 maintains an index $j\in[n]$ throughout its execution such that $c_{j}$ is a candidate $(1+\epsilon)$-approximate nearest neighbor of $q$. Initially, $j$ is set to $1$. The algorithm runs the 2-point comparator $n-1$ times, where in the $i$-th iteration, a comparison is made between $c_{i}$ and the current nearest neighbor candidate $c_{j}$. If $A_{c_{i},c_{j}}$ outputs $c_{i}$, then we know, by Lemma 4.3, that $c_{i}$ is strictly closer to $q$ than $c_{j}$. Additionally, if $c_{i}$ were closer to $q$ by a factor of $1+\epsilon$ or more, then we are guaranteed that $c_{i}$ is output by $A_{c_{i},c_{j}}$.
@@
Since the data structure consists of $n^{2}$ copies of data structure in Algorithm 1, the space complexity of Algorithm 2 is $n^{2}\times O(\log(n^{2}/\delta)/\epsilon^{2})=O(n^{2}\log(n/\delta)\log(d)/\epsilon^{2})$. We make $n-1$ queries to the data structure in Algorithm 1, leading to the query complexity $O(n\log(n/\delta)/\epsilon^{2})$. ∎
5 Near-linear Space Data Structures for $\ell_{1}$ and $\ell_{2}$ Metrics
Theorem 5.1.
We prove the $\ell_{1}$ and $\ell_{2}$ metric cases of Theorem 5.1 separately in two subsections. Both proofs follow the same structure although they differ in details. We consider the following four events and prove that each of them holds with probability $41-\delta/$ 4. We establish the correctness of the data structure by conditioning on the first three events. By conditioning on the last event, we demonstrate the space and query complexity.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
We also show that if all these four events hold, the requirements of Theorem 5.1 are met. Then we use union bound to conclude the theorem.
The preprocessing and querying procedures of our data structure are described in Algorithm 3. The algorithm receives a set $C$ of $n$ points and parameters $\epsilon,\delta\in(0,1)$ as inputs for its preprocessing phase. During preprocessing, it iteratively samples coordinates from $[d]$ to form a multiset $I$  [^4] Multiple coordinates may be added to $I$ in the same iteration, and $I$ may contain repeated coordinates.. Let $R$ denote the set $C$ of $n$ points restricted to $I$. An $\ell_{1}$ dimension reduction is then applied to $R$ to further reduce the space. In the query phase, the algorithm receives a query $q\in\mathbb{R}^{d}$. It probes $q$ on all coordinates in $I$ and applies the same $\ell_{1}$ dimension reduction to the coordinate-selected query. The output, identifying which point in $C$ is the nearest neighbor of $q$, is determined by which point, after coordinate selection and dimension reduction, is closest to the query.
3
4      for $t\in[T]$ do
5             for $b\in[d]$ do
6                   Add $b$ to $I$ with probability $p^{(b)}\triangleq\max_{(i,j)\in{n\choose 2}}\frac{\left|c_{i}^{(b)}-c_{j}^{(b)}\right|}{\|c_{i}-c_{j}\|_{1}}$;
7
8
10       Let $m\leftarrow O(\log(n/\delta)/(\epsilon^{2}\delta^{2}))$;
11       Sample $M\in\mathbb{R}^{[m]\times I}$, where each entry of $M$ follows the Cauchy distribution, whose density function is $c(x)=\frac{1}{\pi(1+x^{2})}$. Notice that $M(\cdot)\colon\mathbb{R}^{I}\rightarrow\mathbb{R}^{m}$ is an oblivious linear mapping [Ind06];
12
13      store $I$, $M(R)=\{Mr_{i}|i\in[n]\}$, $\{p^{(b)}|b\in I\}$, $M$
14Query $(q)$ // Inputs: $q\in\mathbb{R}^{d}$
15       Query $q^{(b)},\forall b\in I$;
16       Let $u\in\mathbb{R}^{I}$, where $\forall b\in I,u^{(b)}=q^{(b)}/p^{(b)}$;
18       Let $\hat{i}=\operatorname*{arg\,min}_{i\in[n]}F(Mr_{i},Mu)$;
19       return $\hat{i}$
Theorem 5.2.
We show the proof in the following subsections in line with the structure described above. Let $\delta^{\prime}=\delta/4$, where each good event described above happens with probability $1-\delta^{\prime}$.
Claim 5.3.
$\frac{1}{T}\|r_{i^{*}}-u\|_{1}\leq\frac{1}{\delta^{\prime}}\|c_{i^{*}}-q\|_{1},$
with probability $1-\delta^{\prime}$.
Proof.

### v2

@@ 3 Preliminaries
4 Near-linear Space Data Structures for $\ell_{1}$ and $\ell_{2}$ Metrics
Theorem 4.1.
The proof of Theorem 4.1 for the case of $\ell_{1}$-metric is presented in Section 4.1 with missing proofs deferred to Section B.1, where as the proof for the case of $\ell_{2}$-metric is in Section B.2. Both proofs follow the same structure although they differ in details.
The preprocessing and querying procedures of our data structure are described in Algorithm 1. The algorithm receives a set $C$ of $n$ points and parameters $\epsilon,\delta\in(0,1)$ as inputs for its preprocessing phase. During preprocessing, it iteratively samples coordinates from $[d]$ to form a multiset $I$ [^5] Multiple coordinates may be added to $I$ in the same iteration, and $I$ may contain repeated coordinates. Let $R$ denote the set $C$ of $n$ points restricted to $I$. An $\ell_{1}$ dimension reduction is then applied to $R$ to further reduce the space. In the query phase, the algorithm receives a query $q\in\mathbb{R}^{d}$. It probes $q$ on all coordinates in $I$ and applies the same $\ell_{1}$ dimension reduction to the coordinate-selected query. The output, identifying which point in $C$ is the nearest neighbor of $q$, is determined by which point, after coordinate selection and dimension reduction, is closest to the query.
3
4      for $t\in[T]$ do
5             for $b\in[d]$ do
6                   Add $b$ to $I$ with probability $p^{(b)}\triangleq\max_{(i,j)\in{\binom{n}{2}}}\frac{\left|c_{i}^{(b)}-c_{j}^{(b)}\right|}{\|c_{i}-c_{j}\|_{1}}$;
7
8
10       Let $m\leftarrow O(\log(n/\delta)/(\epsilon^{2}\delta^{2}))$;
11       Sample $M\in\mathbb{R}^{[m]\times I}$, where each entry of $M$ follows the Cauchy distribution, whose density function is $c(x)=\frac{1}{\pi(1+x^{2})}$. Notice that $M(\cdot)\colon\mathbb{R}^{I}\rightarrow\mathbb{R}^{m}$ is an oblivious linear mapping [23];
12
13      store $I$, $M(R)=\{Mr_{i}|i\in[n]\}$, $\{p^{(b)}|b\in I\}$, $M$
14Query $(q)$ // Inputs: $q\in\mathbb{R}^{d}$
15       Query $q^{(b)},\forall b\in I$;
16       Let $u\in\mathbb{R}^{I}$, where $\forall b\in I,u^{(b)}=q^{(b)}/p^{(b)}$;
18       Let $\hat{i}=\operatorname*{arg\,min}_{i\in[n]}F(Mr_{i},Mu)$;
19       return $\hat{i}$
Theorem 4.2.
To prove Theorem 4.2, we consider the following four events and prove that each of them holds with probability $1-\delta^{\prime}$, where $\delta^{\prime}=\delta/4$. We establish the correctness of the data structure by conditioning on the first three events. By conditioning on the last event, we demonstrate the space and query complexity.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Finally, by applying a union bound, we show that all four events hold with probability $1-\delta$, which implies that the requirements of Theorem 4.1 are met. We show the proof in the following subsections in line with the structure described above.
Claim 4.
Let $q\in\mathbb{R}^{n}$ be a query and let $i^{*}\in[n]$ be the index of its nearest neighbor. It holds that $\frac{1}{T}\|r_{i^{*}}-u\|_{1}\leq\frac{1}{\delta^{\prime}}\|c_{i^{*}}-q\|_{1},$ with probability $1-\delta^{\prime}$.
4.1.2 Coordinate Selection Preserves Distance Ratio
Definition 4.3 (Good region and bad region).
Consider two points $c_{1},c_{2}\in\mathbb{R}^{d}$, and a query point $q\in\mathbb{R}^{d}$. We define the good region $\mathsf{Good}(c_{1},c_{2},q)$ w.r.t. $(c_{1},c_{2},q)$ as
$\{z\in[d]\mid\min\{c_{1}^{(z)},c_{2}^{(z)}\}-\frac{8}{\epsilon\delta^{\prime}}|c_{1}^{(z)}-c_{2}^{(z)}|\leq q^{(z)}\leq\max\{c_{1}^{(z)},c_{2}^{(z)}\}+\frac{8}{\epsilon\delta^{\prime}}|c_{1}^{(z)}-c_{2}^{(z)}|\}.$
Definition 4.4 (Restricted distance).
Lemma 4.5.
Lemma 4.6.
  1. 1.
  2. 2.
  3. 3.
Lemma 4.7.
4.1.3 $\ell_{1}$ Dimension Reduction Preserves the Distance Ratio
Theorem 4.8 (Theorem 1 in [23]).
Claim 5.
$\frac{\|c_{i}-q\|_{1}}{\|c_{i^{*}}-q\|_{1}}\leq 1+\epsilon$
with probability $1-3\delta^{\prime}$.
Proof 4.9.
Applying Theorem 4.8 with $m=O(\log(n/\delta^{\prime})/(\epsilon^{2}\delta^{\prime 2}))$, given any $v\in R$, it holds with probability $1-\delta^{\prime}/(n+1)$, that $F(Mv,Mu)\in[1-\epsilon\delta^{\prime}/100,1+\epsilon\delta^{\prime}/100]\|v-u\|_{1}$. By union bound, with probability $1-\delta^{\prime}$, it holds for all $v\in R$. By Lemma 4.7 it holds with probability $1-2\delta^{\prime}$ that for all $i\in[n]$, if $\frac{\|c_{i}-q\|_{1}}{\|c_{i^{*}}-q\|_{1}}\geq 1+\epsilon$, then $\frac{\|r_{i}-u\|_{1}}{\|r_{i^{*}}-u\|_{1}}\geq 1+\epsilon\delta^{\prime}/4$, thus $\frac{F(Mr_{i},Mu)}{F(Mr_{i^{*}},Mu)}>1$ with probability $1-3\delta^{\prime}$, i.e. the algorithm will not output $i$ as the index of the approximate nearest neighbor.
4.1.4 Space and Query Complexity
Lemma 4.10.
$1\leq\sum_{b\in[d]}p^{(b)}\leq n$.
Proof 4.11.
Claim 6.
Proof 4.12.
Notice that $f(x)$ is a non-decreasing function, thus $f(x)\geq f(y)$, which proves $x-y\geq x^{\prime}-y^{\prime}$. Notice that $0\leq f(x)\leq\tau_{2}-\tau_{1}$, thus $f(x)\leq f(y)+(\tau_{2}-\tau_{1})$, which proves $x^{\prime}-y^{\prime}\geq x-y-(\tau_{2}-\tau_{1})$.
Claim 7.
Proof 4.13.
To prove (ii), define $p_{b}=\max_{c_{i},c_{j}\in C,i\neq j}\frac{|c_{i}^{(b)}-c_{j}^{(b)}|}{\|c_{i}-c_{j}\|_{1}}$ and $p_{b}^{\prime}=\max_{c^{\prime}_{i},c^{\prime}_{j}\in C^{\prime},i\neq j}\frac{|{c^{\prime}_{i}}^{(b)}-{c^{\prime}_{j}}^{(b)}|}{\|c^{\prime}_{i}-c^{\prime}_{j}\|_{1}}$. We prove that $p_{b}-p_{b}^{\prime}\leq\frac{|c_{i_{0}}^{b}-c_{j_{0}}^{b}|}{\|c_{i_{0}}-c_{j_{0}}\|_{1}},\forall b\in[d]$. The claim follows, since $P_{k}(C)-P_{k-1}(C^{\prime})=\sum_{b}(p_{b}-p_{b}^{\prime})\leq\sum_{b}\frac{|c_{i_{0}}^{(b)}-c_{j_{0}}^{(b)}|}{\|c_{i_{0}}-c_{j_{0}}\|_{1}}=1$.
If ${c^{\prime}_{u}}^{(b)}\neq{c^{\prime}_{v}}^{(b)}$, ${p_{b}}^{\prime}\geq\frac{|{c^{\prime}_{u}}^{(b)}-{c^{\prime}_{v}}^{(b)}|}{\|c^{\prime}_{u}-c^{\prime}_{v}\|_{1}}\geq\frac{|{c_{u}^{(b)}}-{c_{v}^{(b)}}|-|{c_{i_{0}}^{(b)}}-{c_{j_{0}}^{(b)}}|}{\|c_{u}-c_{v}\|_{1}}.$ The second inequality follows from 6. Thus $p_{b}-p_{b}^{\prime}\leq\frac{|{c_{i_{0}}^{(b)}}-{c_{j_{0}}^{(b)}}|}{\|c_{u}-c_{v}\|_{1}}\leq\frac{|{c_{i_{0}}^{(b)}}-{c_{j_{0}}^{(b)}}|}{\|c_{i_{0}}-c_{j_{0}}\|_{1}}.$ If ${c^{\prime}_{u}}^{(b)}={c^{\prime}_{v}}^{(b)}$, it means $|{c_{u}^{(b)}}-{c_{v}^{(b)}}|\leq|{c_{i_{0}}^{(b)}}-{c_{j_{0}}^{(b)}}|$, thus $p_{b}-p_{b}^{\prime}\leq p_{b}=\frac{|c_{u}^{(b)}-c_{v}^{(b)}|}{\|c_{u}-c_{v}\|_{1}}\leq\frac{|{c_{i_{0}}^{(b)}}-{c_{j_{0}}^{(b)}}|}{\|c_{i_{0}}-c_{j_{0}}\|_{1}}.$
Solving the recurrence given by 7 with the boundary condition that for a set $C^{\prime\prime}$ of two points, $P_{2}(C^{\prime\prime})=1$, we get that $P_{n}(C)\leq n$. Since $P_{n}(C)$, for the set of $n$ points, is by definition, equal to $\sum_{b\in[d]}p^{(b)}$, the statement of the lemma follows.
Claim 8.
With probability $1-\delta^{\prime}$, we have $|I|\leq 2Tn$.
Proof 4.14.
We define $X_{t,z}$ for $t\in[T],z\in[d]$ to be the indicator random variable for the event that we add $z$ to $I$ at iteration $t$. Note that $\Pr[X_{t,z}=1]=p^{(z)}$. We can see that $|I|=\sum_{t\in[T],z\in[d]}X_{t,z}$. Hence, $\mathbb{E}[|I|]=\sum_{t\in[T],z\in[d]}p^{(z)}\leq T\cdot\sum_{b\in[d]}p^{(b)}\leq T\cdot n$, where the last inequality follows from Lemma 4.10. Similarly, we can see that $\mathbb{E}[|I|]\geq T$. Using Chernoff bound, we have $\Pr\left[|I|-\mathbb{E}\left[|I|\right]\geq\mathbb{E}\left[|I|\right]\right]\leq\exp(-\mathbb{E}[|I|]/3)\leq\exp(-T/3)\leq\delta/4.$ Thus, with probability $41-\delta/$ 4, we have $|I|\leq 2E[|I|]\leq 2Tn$.
Lemma 4.15.
With prob. $1-\delta^{\prime}$, the space complexity of Algorithm 1 is $O(\frac{n\log^{2}(n/\delta)\log(d)}{\epsilon^{5}\delta^{4}})$ wordsize, and the number of coordinates that we query is $O(\frac{n\log(n/\delta)}{\epsilon^{3}\delta^{2}})$.
Proof 4.16.
By 8, we have that the number of coordinates of $q$ that we query is $|I|=O(Tn)=O(\frac{n\log(n/\delta)}{\epsilon^{3}\delta^{2}})$. Next we analyze the space complexity of the data structure. Matrix $M$ is of wordsize $O(m|I|)$, $M(R)$ is of wordsize $O(mn)$, $I$ is of wordsize $O(|I|\log(d))$. Thus the space complexity of the our data strcuture is $O(m|I|+mn+|I|\log(d))=O(\frac{n\log^{2}(n/\delta)\log(d)}{\epsilon^{5}\delta^{4}})$ wordsize.
@@ 4.1.4 Space and Query Complexity
Appendix A Warm-Up: Data Structure for $\ell_{1}$ Metric with $\tilde{O}(n^{2})$ Space and $\tilde{O}(n)$ Query Complexity
@@
Definition A.1 (Scaled Bounding Box).
For any two points $a,b\in\mathbb{R}^{d}$, the scaled bounding box $B_{a,b}$ specified by $(a,b)$ is
$\left\{x\in\mathbb{R}^{d}\mid\forall z\in[d],\min\{a^{(z)},b^{(z)}\}-100\cdot|a^{(z)}-b^{(z)}|\leq x^{(z)}\right.$
$\left.\leq\max\{a^{(z)},b^{(z)}\}+100\cdot|a^{(z)}-b^{(z)}|\right\}.$
@@
Theorem A.2 is a special case of Theorem D.1, where a rigorous proof is provided. In this section, we present a more concise proof to provide intuition. Our approach is the following. We first describe a data structure for the approximate nearest neighbor problem for two points (i.e. a 2-point comparator), and then employ a naive min-finding procedure with oracle access to the 2-point comparator, to extend to the approximate nearest neighbor with $n$ points.
A.1 A Strong Comparator for Two Points
@@
We analyze the correctness of Algorithm 2 in Lemma A.3. In particular, we show the following strong guarantees: (i) if one point is significantly closer to the query than the other point, the comparator outputs the correct point with probability $1-\delta$, (ii) otherwise, the comparator may output a special symbol $\perp$, but never outputs a wrong point as the nearest neighbor.
@@ 10       Let $x,y\leftarrow 0$;
Lemma A.3 (Strong Comparator Correctness).
@@
Since $\mathbb{E}[Y_{t}]=\|b-q\|_{1}\geq\frac{1}{2}\|b-a\|_{1}$, using Chernoff bound, we have $\Pr\left[|y-\mathbb{E}[y]|>\frac{\epsilon}{8}\mathbb{E}[y]\right]\allowbreak\leq\frac{\delta}{2}$ by choosing $T=\Theta(\log(1/\delta)/\epsilon^{2})$. To analyze $x$, we consider the following two cases:
Case 1. If $\mathbb{E}[X_{t}]=\|a-q\|_{1}\geq\frac{1}{8}\|b-a\|_{1}$, then using Chernoff bound, we have
$\Pr\left[|x-\mathbb{E}[x]|>\frac{\epsilon}{8}\mathbb{E}[x]\right]\leq\frac{\delta}{2}$ by choosing $T=\Theta(\log(1/\delta)/\epsilon^{2})$.
Case 2. If $\mathbb{E}[X_{t}]=\|a-q\|_{1}<\frac{1}{8}\|b-a\|_{1}$, then using Chernoff bound, we have
$\Pr\left[|x-\mathbb{E}[x]|>\frac{T}{8}\|b-a\|_{1}\right]\leq\frac{\delta}{2}$ by choosing $T=\Omega(\log(1/\delta))$.
To prove item 1 of Lemma A.3, suppose that $\frac{\|b-q\|_{1}}{\|a-q\|_{1}}\geq 1+\epsilon$. If we are in Case 1, with probability $1-\delta$ we have $x\approx_{\frac{\epsilon}{8}}T\|a-q\|_{1}$ and $y\approx_{\frac{\epsilon}{8}}T\|b-q\|_{1}$. This implies that, with probability $1-\delta$ we get $\frac{y}{x}\geq\frac{1-\frac{\epsilon}{8}}{1+\frac{\epsilon}{8}}\cdot(1+\epsilon)\geq 1+\frac{\epsilon}{2}$. Thus the querying phase correctly outputs $a$ as the nearest neighbor. If we are in Case 2, with probability $1-\delta$ we have $x\leq\frac{1}{4}T\|b-a\|_{1}$, and $y\approx_{\frac{\epsilon}{8}}T\|b-q\|_{1}$, which implies that $\frac{y}{x}\geq\frac{(1-\frac{\epsilon}{8})(1-\frac{1}{8})}{\frac{1}{4}}\geq 3$. Thus the querying phase correctly outputs $a$.
The second item of Lemma A.3 can be proved similarly.
A.2 Data Structure for $\ell_{1}$ Metric with $\tilde{O}(n^{2})$ Space
@@ 10                   Let $j\leftarrow i$;
Proof A.5.
(for Theorem A.2) By a union bound, we get that with probability $1-\delta$, for all $1\leq i<j\leq n$, the data structure $A_{(c_{i},c_{j})}$ does not fail. We condition on this event in the rest of the proof. Notice that $j$ is a variable under updating.
Algorithm 3 maintains an index $j\in[n]$ throughout its execution such that $c_{j}$ is a candidate $(1+\epsilon)$-approximate nearest neighbor of $q$. Initially, $j$ is set to $1$. The algorithm runs the 2-point comparator $n-1$ times, where in the $i$-th iteration, a comparison is made between $c_{i}$ and the current nearest neighbor candidate $c_{j}$. If $A_{c_{i},c_{j}}$ outputs $c_{i}$, then we know, by Lemma A.3, that $c_{i}$ is strictly closer to $q$ than $c_{j}$. Additionally, if $c_{i}$ were closer to $q$ by a factor of $1+\epsilon$ or more, then we are guaranteed that $c_{i}$ is output by $A_{c_{i},c_{j}}$.
@@
Since the data structure consists of $n^{2}$ copies of data structure in Algorithm 2, the space complexity of Algorithm 3 is $n^{2}\times O(\log(n^{2}/\delta)/\epsilon^{2})=O(n^{2}\log(n/\delta)\log(d)/\epsilon^{2})$. We make $n-1$ queries to the data structure in Algorithm 2, leading to the query complexity $O(n\log(n/\delta)/\epsilon^{2})$.
Appendix B Missing Proofs from Section 4 for $\ell_{1}$-metric and complete proof $\ell_{2}$-metric
B.1 Missing Proofs from Section 4 for $\ell_{1}$-metric
Proof B.1 (Proof of 4).